OpenStudy (anonymous):

Let g(x) = -x2 + x. Find g(-10). A) -110 B) -90 C) 0 D) 90 Eliminate

OpenStudy (raden):

just put x=-10

OpenStudy (raden):

g(x) = -x2 + x g(-10) = -(-10)^2 + (-10) = -100 - 10 = ...

OpenStudy (anonymous):

Thanks that helped a lot.

OpenStudy (raden):

yw